397 C 3 =10,349,790
How do you get this answer?

Answer by sudhanshu_kmr(1152)

nCr = n!/ [(n-r)! * r!]

n! = n * (n-1)* (n-2) *.....*1

example.. 5! = 5*4*3*2*1 = 120

397 C 3 = 397! / [ 394! * 3!]